Question 1
-The 112 students sat for an examination.Question 1 was answered by 50 students,Question 2 by 66 students,Question 3 by 38 students.32 students attempeted Q2 and Q1,22 students attempted Q2&Q3,20 students attempted Q3&Q1,8 students attempted all the questions.How many students didn't attempt any questi
Question 2
Supposed 2 coins are tosed,events A and B are defined as follows; A is the event of getting a head in the first coin.B is the event that the coin falls alike.Fknd the probability of the event A happening given that B has happened and the probability of event B happening given that A has already happened.
Question 3
Past records shown that the probability of a school football twam winning a match is 0.7.How many possible games might the school win in 30 matches

Answer by ikleyn(52781)

.
Question 1
The 112 students sat for an examination.
Question 1 was answered by 50 students, Question 2 by 66 students, Question 3 by 38 students.
32 students attempeted Q2 and Q1, 22 students attempted Q2 and Q3, 20 students attempted Q3 Q1,
8 students attempted all the questions. How many students didn't attempt any questi
Question 2
Supposed 2 coins are tosed,events A and B are defined as follows; A is the event of getting a head in the first coin.
B is the event that the coin falls alike.Fknd the probability of the event A happening given that B has happened
and the probability of event B happening given that A has already happened.
Question 3
Past records shown that the probability of a school football twam winning a match is 0.7.
How many possible games might the school win in 30 matches
~~~~~~~~~~~~~~~~~~~

            I will answer Q1 and Q3, only.


                    Question 1

Use the Inclusion-Exclusion principle (formula).  The number of those who attempted at least one of the questions is

    n(Q1 U Q2 U Q3) = n(Q1) + n(Q2) + n(Q3) - n(Q1 and Q2) - n(Q1 and Q3) - n(Q2 and Q3) + n(Q1 and Q2 and Q3) = 

                    =  50   +  66   +  38   -     32       -     20       -     22       +          8          = 88.

Thus the number of those who didn't attempt any question is the COPMPLEMENT of 88 to 112, i.e.  112 - 88 = 24.    ANSWER


                    Question 3

They ask about the mathematical expectation.  You may use your intuition (= common sense) or strict mathematical formula

for the mathematical expectation of the binomial distribution with 30 trials and the probability of success p = 0.7 
for each individual trial.


In any case, the ANSWER  is  this number  E = n*p = 0.7*30 = 21.

Solved.

------------------

On inclusion-exclusion principle,  see this Wikipedia article

https://en.wikipedia.org/wiki/Inclusion%E2%80%93exclusion_principle


To see many other similar  (and different)  solved problems on Inclusion-Exclusion,  see the lessons

    - Counting elements in sub-sets of a given finite set
    - Advanced problems on counting elements in sub-sets of a given finite set
    - Challenging problems on counting elements in subsets of a given finite set
    - Selected problems on counting elements in subsets of a given finite set
    - Inclusion-Exclusion principle problems

in this site.


Happy learning (!)